### v1

# A Survey on LLM-as-a-Judge

Abstract.
Abstract
Accurate and consistent evaluation is crucial for decision-making across numerous fields, yet it remains a challenging task due to inherent subjectivity, variability, and scale. Large Language Models (LLMs) have achieved remarkable success across diverse domains, leading to the emergence of ”LLM-as-a-Judge,” where LLMs are employed as evaluators for complex tasks. With their ability to process diverse data types and provide scalable, cost-effective, and consistent assessments, LLMs present a compelling alternative to traditional expert-driven evaluations. However, ensuring the reliability of LLM-as-a-Judge systems remains a significant challenge that requires careful design and standardization. This paper provides a comprehensive survey of LLM-as-a-Judge, addressing the core question: How can reliable LLM-as-a-Judge systems be built? We explore strategies to enhance reliability, including improving consistency, mitigating biases, and adapting to diverse assessment scenarios. Additionally, we propose methodologies for evaluating the reliability of LLM-as-a-Judge systems, supported by a novel benchmark designed for this purpose. To advance the development and real-world deployment of LLM-as-a-Judge systems, we also discussed practical applications, challenges, and future directions. This survey serves as a foundational reference for researchers and practitioners in this rapidly evolving field. https://github.com/IDEA-FinAI/LLM-as-Evaluator.
copyright: none
1. Introduction
Judgment is the faculty of thinking the particular as contained under the universal. It involves the capacity to subsume under rules, that is, to distinguish whether something falls under a given rule. —— Kant, Critique of Judgment (Kant, 1790), Introduction IV, 5:179; Critique of Pure Reason (Kant, 1781), A132/B171.
Recently, Large Language Models (LLMs) have achieved remarkable success in numerous domains, ranging from artificial intelligence and software engineering to education and social science. The adoption of LLMs as evaluators—commonly referred to as ”LLM-as-a-Judge” (Zheng et al., 2023b)—has surged, driven by their ability to emulate human-like reasoning and decision-making processes. This capability enables LLMs to undertake roles traditionally reserved for human experts, offering a cost-effective and scalable alternative. For instance, we usually rely on experts to evaluate the accuracy of mathematics and physics competition questions at the Olympiad level (He et al., 2024a), which can be assessed through LLM-as-a-Judge now. Additionally, in recent peer reviews of research submissions, LLM-as-a-Judge is introduced to address the rising number of paper submissions and reviewer workload [^1] https://blog.iclr.cc/2024/10/09/iclr2025-assisting-reviewers/, which is designed to identify potential issues in reviews and offer constructive feedback to reviewers. These trends underscore a key motivation for adopting LLM-as-a-Judge: the potential to enhance evaluation efficiency while addressing limitations inherent in human assessments, such as scalability and consistency.
LLM-as-a-Judge presents a compelling alternative to both human evaluations and traditional automated methods, offering distinct advantages in scalability, efficiency, and adaptability. Human evaluations, though often considered the gold standard, face challenges in scalability, cost, and consistency. They are time-consuming, require substantial expert effort, and are expensive to scale due to limited availability of qualified evaluators. Coordinating and training evaluators adds complexity, and fatigue during lengthy tasks can compromise reliability and accuracy. In contrast, LLMs provide scalable, cost-effective, and efficient evaluations with reduced subjective variability, enhancing objectivity. Traditional automated methods, such as BLEU, ROUGE, and METEOR for software artifact summarization, often fail to align with human judgment or provide clear insights in specialized domains like software engineering. LLMs offer flexibility to process diverse input types, including text, semi-structured data, and multi-modal content, allowing evaluations to integrate qualitative insights with quantitative rigor. This human-aligned adaptability makes LLMs effective for complex, context-aware assessments beyond the limits of conventional metrics (Wang et al., 2023d; Li et al., 2023b; Chen et al., 2024c).
Despite its wide advantages, LLM-as-a-Judge poses significant challenges for reliability. This necessitates the capability of LLM-as-a-Judge framework to subsume under rules, that is, to distinguish whether something falls under a given rule (Kant, 1781). As LLM-as-a-Judge becomes more commonly used as an effective evaluator in different areas, collecting evaluations with LLM-as-a-Judge is relatively simple. Therefore, central to this survey is the fundamental question: How to build reliable LLM-as-a-Judge systems? To address this question, we explore two core aspects: (1) strategies for enhancing the reliability of LLM-as-a-Judge and (2) methodologies for evaluating reliability of LLM-as-a-Judge systems themselves.
For the first aspect of enhancing LLM-as-a-Judge reliability, we review the main strategies aimed at optimizing their performance for diverse evaluation tasks. These strategies include improving consistency, mitigating biases, and refining adaptability to different assessment scenarios. For the second aspect, we examine the metrics, datasets, and methodologies used to evaluate the performance of LLM-as-a-Judge systems, discussing potential sources of bias and corresponding mitigation techniques. Building on this foundation, we introduce a novel benchmark specifically designed for LLM-as-a-Judge evaluations. Using established metrics and datasets, this benchmark provides a framework for analyzing the effectiveness of various reliability enhancement strategies. Additionally, we explore practical application scenarios, identify specific challenges unique to each context, and propose solutions to address these issues. Finally, we discuss future research directions, emphasizing key areas for advancing the reliability, scalability, and applicability of LLM-as-a-Judge systems.
This paper aims to provide a comprehensive overview of the LLM-as-a-Judge research landscape while offering insights into how reliable LLM-as-a-Judge can be constructed. We hope that this work will serve as a valuable reference for researchers and practitioners, fostering further research and facilitating the real-world deployment of LLM-as-a-Judge. The rest of this survey is organized as Figure 1. In Section 2, we provide a comprehensive overview of the LLM-as-a-Judge field. We define LLM-as-a-Judge through formal and informal definitions, and categorize existing methods and approaches for its use. For a quick guide on implementing an LLM-as-a-Judge for specific scenarios, you can find answers in Quick Practice (2.5). Then, we discuss the problems of ”How to improve” and ”how to evaluate” LLM-as-a-Judge in n Section 3 to 4 . Next, we show the applications of LLM-as-a-Judge in Section 6. The discussions of the challenges and future directions come at last in Section 7 and Section 8.
Figure 1. Paper Structure.
2. Background and Method
In evaluative tasks, especially those that are subjective, human assessment is often considered the gold standard due to its reliable and open-ended nature (Shi et al., 2024; Gao et al., 2023). However, this approach is typically slow and costly (Zheng et al., 2023b; Zhu et al., 2023a; Yuan et al., 2023). To address these challenges, LLMs are increasingly employed as substitutes for human evaluators. Since these models are frequently trained using Reinforcement Learning from Human Feedback (RLHF), they demonstrate strong alignment with human perspectives, leading to the approach known as ”LLM-as-a-Judge”.
In general, LLM-as-a-Judge is a process that uses LLM to evaluate different objects in various scenarios for diverse tasks. For instance, roles such as ”Assessors”, ”Critics,” and ”Verifiers” utilize LLMs to facilitate evaluation at different stages of the process, whether during intermediate steps or throughout the entire workflow. To date, definition of how to effectively use LLM-as-a-Judge for evaluation tasks has been largely informal or vague, lacking clear and formal expression. Therefore, we will provide a formal definition of LLM-as-Evaluator as follows:
[…]
  • $\mathcal{E}$: The final evaluation obtained from the whole LLM-as-a-Judge process in the expected manner. It could be a score, a choice, or a sentence, etc.
  • $\mathcal{P}_{\mathcal{LLM}}$: The probability function defined by the corresponding LLM, and the generation is an auto-regressive process.
  • $x$: The input data in any available types (text, image, video), which waiting to be evaluated.
  • $\mathcal{C}$: The context for the input $x$, which is often prompt template or combined with history information in dialogue.
  • $\oplus$: The combination operator combines the input $x$ with the context $\mathcal{C}$, and this operation can vary depending on the context, such as being placed at the beginning, middle, or end.
The formulation of LLM-as-a-Judge reflects that LLM is a type of auto-regressive generative model, which generates subsequent content based on the context then obtain target evaluation from it. The form of LLM-as-a-Judge illustrates how we utilize LLM for evaluation tasks, encompassing input design, model selection and training, as well as output post-processing. The different basic approaches of implementing LLM-as-a-Judge can be classified according to the formulation: In-Context Learning, Model Selection, Post-processing Method and Evaluation Pipeline, which concluded in Figure 2. By following this pipeline, we can build a basic LLM-as-a-Judge for evaluation. A faster practice guide is available in section 2.5.
Figure 2. LLM-as-a-Judge Evaluation Pipelines.
2.1. In-Context Learning
To apply LLM-as-a-Judge, it is helpful to start by defining the evaluation task using In-Context Learning methods. This process involves two key aspects: the design of prompt and input. For input design, it is important to consider the type of variables to be evaluated (such as text, image, or video), the manner of input (e.g., individually, in pairs, or in batches), and the position of the input (e.g., at the beginning, middle, or end). As for the prompt design, four different methods can be adopted, as illustrated in Figure 2. The four methods include generating scores, solving true/false questions, conducting pairwise comparisons, and making multiple-choice selections. Further details will be provided in the following sections.
2.1.1. Generating scores
It is quite intuitive to represent an evaluation using a corresponding score. What requires more careful consideration, however, is the nature and range of the score used for evaluation. The score can be discrete, with common ranges like 1-3, 1-5 (Jones et al., 2024), or 1-10 (Zhu et al., 2023a; Li et al., 2023b). Alternatively, it can be continuous, ranging from 0 to 1 or 0 to 100 (Xiong et al., 2024). The simplest way to score is through the context, setting the range of scores and the main criteria for scoring. For example, ”Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance” (Zhu et al., 2023a). A slightly more complex way is to provide more detailed scoring criteria. More complex scoring situations can be as Language-Model-as-an-Examiner (Bai et al., 2024a), which use Likert scale scoring functions as an absolute evaluative measure showed in Figure 3. The evaluator assigns scores to a given response along predefined dimensions including accuracy, coherence, factuality and comprehensiveness. Each of these dimensions is scored on a scale of 1 to 3, ranging from worst to best. The evaluator is also asked to provide an overall score ranging from 1 to 5, based on the scores assigned to the previous 4 dimensions. This score serves as an indicator of the overall quality of the answer.
Evaluate the quality of summaries written for a news article. Rate each summary on four dimensions: {Dimension_1}, {Dimension_2}, {Dimension_3}, and {Dimension_4}. You should rate on a scale from 1 (worst) to 5 (best). Article: {Article} Summary: {Summary}
Figure 3. The template for Likert scale scoring from Gao et al. (2023).
2.1.2. Solving Yes/No questions
A Yes/No question requires a judgment on a given statement, focusing solely on its accuracy. This type of question is simple and direct, providing only two fixed responses—yes or no, true or false—without any additional comparisons or choices.
This type of evaluation is often utilized in intermediate processes, creating the conditions for a feedback loop. For example, it promotes a self-optimization cycle, as seen in Reflexion (Shinn et al., 2024), which generates verbal self-reflections to provide valuable feedback for future attempts. In scenarios with sparse reward signals, such as a binary success status (success/fail), the self-reflection model uses the current trajectory and persistent memory to generate nuanced and specific feedback. Similarly, in self-improvement contexts (Tian et al., 2023b), Yes/No questions can be employed to evaluate custom phrases, such as "Modification needed." and "No modification needed.", facilitating entry into the next cycle. Moreover, these evaluations are common for testing knowledge accuracy and assessing whether statements align with established facts (Sun et al., 2023c), like ”Given a question and the associated retrieved knowledge graph triples (entity, relation, entity), you are asked to answer whether it’s sufficient for you to answer the question with these triples and your knowledge (Yes or No).” A detailed and specific example can be seen in the Figure 4.
Is the sentence supported by the article? Answer ”Yes” or ”No”. Article: {Article} Sentence: {Sentence}
Figure 4. The template for Yes/No evaluation for example.
2.1.3. Conducting pairwise comparisons
Pairwise comparison refers to comparing two options and selecting which one is superior or more aligned with a specific standard, showed in Figure 5. It involves making a decision between two options rather than judgement between ’yes’ or ’no’. The comparison can be subjective or based on objective criteria. This evaluation is a relative evaluation. Pairwise comparison is often used for ranking multiple options or prioritizing them, where several comparisons are made between pairs to identify the better choice or establish a hierarchy.
Pairwise comparison is a well-established method that has significantly impacted a variety of fields (Qin et al., 2023). As noted by (Liu et al., 2024c), LLM and human evaluations are more aligned in the context of pairwise comparisons compared to score-based assessments. Numerous studies have demonstrated that pairwise comparative assessments outperform other judging methods in terms of positional consistency (Zheng et al., 2023c; Liusie et al., 2024). Furthermore, pairwise comparisons can be extended to more complex relation-based assessment frameworks, such as list-wise comparisons, using advanced ranking algorithms (Qin et al., 2023; Liu et al., 2024c), data filtering (Yuan et al., 2023). In pairwise comparative assessments, LLM-as-a-Judge is prompted to select the response that better answers the question at hand. To accommodate the possibility of a tie, several option modes are introduced. The Two-Option mode requires judges to choose the better response from two given options. The Three-Option mode introduces an additional choice, allowing judges to indicate a tie if neither response is preferable. Evaluations typically involve determining the outcomes of win, tie, or loss for responses (Wang et al., 2023d) through pairwise comparisons, with win rounds counted for each response. The Four-Option mode further expands the choices, allowing judges to classify responses as either a ”both good tie” or a ”both bad tie.”
Given a new article, which summary is better? Answer ”Summary 0” or ”Summary 1”. You do not need to explain the reason. Article: {Article} Summary 0: {Summary_0} Summary 1: {Summary_1}
Figure 5. The template for pairwise comparison from Gao et al. (2023)
2.1.4. Making multiple-choice selections
Multiple-choice selections involve providing several options, not giving relative choices in pairwise comparison, nor making a yes/no judgment. The evaluator must choose the most appropriate or correct one. This method allows for a broader range of responses compared to true/false questions and can assess deeper understanding or preferences and an example is showed in Figure 6. However, this kind of prompt design is more rare than the first three.
You are given a summary and some semantic content units. For each semantic unit, choose those can be inferred from the summary, return their number. Summary: {Summary} Semantic content units: 1. {SCU_1} 2. {SCU_2} …… n. {SCU_n}
Figure 6. The template for multiple-choice for example.
2.2. Model Selection
2.2.1. General LLM
To automate evaluation by LLM-as-a-Judge, one effective approach is to employ advanced language models such as GPT-4 (OpenAI, 2023a) instead of human evaluators (Zheng et al., 2023c). For instance, Li et al. (2023c) created a test set with 805 questions and assessed the performance by comparing it to text-davinci-003 using GPT-4. Additionally, Zheng et al. (2023c) designed 80 multi-round test questions across eight common areas and used GPT-4 to automatically score the model’s responses. The accuracy of the GPT-4-based evaluator has been demonstrated to be high compared to professional human evaluators, showing superior consistency and stability in evaluations. At the same time, if the general LLM used has limitations in instruction-following or reasoning abilities, the effectiveness of the LLM-as-a-Judge method may be significantly affected.
2.2.2. Fine-tuned LLM
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models specialized in evaluations. For instance, PandaLM (Wang et al., 2023d) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B (Touvron et al., 2023a) as an evaluator model. JudgeLM (Zhu et al., 2023a) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna (Touvron et al., 2023b) as a scalable evaluator model. Auto-J (Li et al., 2023b) constructs evaluation data upon multiple scenarios to train a generative evaluator model, which can provide both evaluation and critical opinion. Prometheus (Kim et al., 2023) defines thousands of evaluation criteria and construct a feedback dataset based on GPT-4, and fine-tunes a fine-grained evaluator model. The typical process for fine-tuning a judge model involves three main steps. Step 1: Data Collection. The training data generally consists of three components: instructions, the objects to be evaluated, and evaluations. Instructions are typically sourced from instruction datasets, while evaluations can come from either GPT-4 or human annotations. Step 2-Prompt Design. The structure of the prompt template can vary based on the evaluation scheme, which already detailed in § 2.1. Step 3: Model Fine-Tuning. Using the designed prompts and collected data, the fine-tuning process for the evaluator model typically adheres to the instruction fine-tuning paradigm (Ouyang et al., 2022b). The model receives an instruction along with one or more responses to generate output that includes evaluation results and possibly explanations.
After fine-tuning, the evaluator model can be employed to evaluate the target object. While these fine-tuned models often demonstrate superior performance on self-designed test sets, they are identified several limitations in their evaluation capabilities, which detailed in Section 4.2.
2.3. Post-processing Method
Post-processing refines the probability distributions generated by LLM-as-a-Judge to provide accurate evaluations. The evaluation format should align with our In-Context Learning design. Additionally, post-processing may involve procedures to enhance the reliability of extracted evaluations, closely linked to the In-Context Learning framework and consistently applied. There are three main methods of post-processing, which are extracting specific tokens, normalizing the output logits, and selecting sentences with high returns.
2.3.1. Extracting specific tokens
As showed in In-context Learning (Section 2.1), when the evaluation target take the form of a score, selecting specific options, or responding with Yes/No, applying rule-match to extract the corresponding token from the response generated during probability distribution iteration is common used. It is worth noting that Yes/No is a broad definition, including custom statements involving judgment. Considering a Yes/No question for evaluation in custom phrases (Tian et al., 2023b): "Modification needed." and "No modification needed." or a yes-no question "Does the above answer need to be further modified?". When the input sample is put through the template, it might have outputs such as ”Modification needed.”, ”Conclusion: Modification needed.” or ”Yes”. This variance in response formats is difficult to parse consistently. The corresponding post-processing with the response is necessary. Using rules to extract specific tokens for our designed prompts and input content, as well as the backbone model used for the evaluator, all have higher requirements as we discussed in Section 2.2. In contextual learning, if there is no clear indication of the output format for response, there may be various expressions of evaluation, which can be seen in Figure 2. For example, ”Response 1 is better” and ”The better one is response 1”, which convey the same choice but differ in format leading to the difficulty of rule recognition. Simple solutions often involve providing clear instructions, such as ”The last sentence should be started with ’The better response is’”, or using a few-shot strategy. Also, the general model with insufficient instruction following capability may not be able to generate the evaluation format and content of the target according to the instruction, resulting in the post-processing extracted according to the rules not as smooth as expected.
2.3.2. Normalizing the output logits
[…]
$\underrightarrow{\overbrace{\left(x\oplus\mathcal{C}\right)}^{\rho_{\text{SC}}}\overbrace{\texttt{"Yes"}}^{\rho_{\text{SR}}}}\ \Rightarrow\begin{cases}\rho_{\text{SC}}=\prod_{t_{i}\in\alpha}P(t_{i}|t_{<i})\cdot\prod_{t_{i}\in\beta}P(t_{i}|t_{<i})\\ \rho_{\text{SR}}=\prod_{t_{i}\in\texttt{"Yes"}}P(t_{i}|t_{<i})\end{cases}$
In addition, Self-evaluation (Hao et al., 2023) is also common using this method for LLM-as-a-Judge. It can be helpful to let the LLM evaluate itself by asking, ”Is this reasoning step correct?” and then reward it based on the probability of the next word being ”Yes.”
2.3.3. Selecting sentences
In addition to selecting specific tokens and normalizing the output logits, the content extracted by LLM-as-a-Judge may also be a sentence or paragraph. As showed in Figure 2, agent for reasoning task (Hao et al., 2023), builds a reasoning tree by iteratively considering the most promising reasoning steps (actions, sub-questions) by LLM-as-a-Judge.
2.4. Evaluation Pipeline
There are three common scenarios for using LLM-as-a-Judge evaluation pipelines showed in Figure 2, which are LLM-as-a-Judge for LLMs, LLM-as-a-Judge for data, and LLM-as-a-Judge for agent respectively.
2.4.1. LLM-as-a-Judge for model
It is universally known that the best way to evaluate LLMs is human judgment, but collecting human annotations can be costly, time-consuming, and laborious (Ouyang et al., 2022b; Zheng et al., 2023d). Using strong LLMs (usually closed-source ones, e.g., GPT-4, Claude, ChatGPT) as an automated proxy for assessing LLMs has become a natural choice (Zhou et al., 2023). With appropriate prompt design, the quality of evaluation and agreement to human judgment can be promising (Dubois et al., 2023; Zheng et al., 2023d; Zhang et al., 2023d; Wang et al., 2023b). However, the cost concern still exists when calling the APIs of these proprietary models, especially when there is a frequent need for model validation on large-scale data. Moreover, closed-source LLM-as-a-Judge leads to low reproducibility due to potential changes in models behind the API. Some recent works have started to make attempts for open-source alternatives. SelFee (Ye et al., 2023) collects generations, feedback, and revised generations from ChatGPT and fine-tunes LLaMA models to build a critique model. Shepherd (Wang et al., 2023c) trains a model that can output critiques for single-response with the data of feedback from online communities and human annotation. PandaLM (Wang et al., 2023d) trains a model to conduct pairwise comparison for LLM Instruction Tuning Optimization, and Zheng et al. (2023d) also fine-tune Vicuna (Touvron et al., 2023b) on a 20K pairwise comparison dataset to explore the potential of open-source models as a more cost-friendly proxy.
Recent advancements in using Large Multimodal Models (LMMs) as evaluators have showcased their potential to perform complex judgment tasks in vision-language scenarios. Proprietary models like GPT-4V and GPT-4o have been pivotal in benchmarks such as detailed captioning and visual chats, utilizing both pointwise and pairwise evaluation methods (Liu et al., 2024a; Zhang et al., 2023b; Lu et al., 2024). Open-source alternatives have emerged, with Prometheus-Vision (Lee et al., 2024a) being the first vision-language model specifically trained to act as an evaluator for user-designed scoring criteria. While Prometheus-Vision introduced the concept of open-source evaluators with a focus on specialized tasks, it remains limited to predefined criteria. In contrast, LLaVA-Critic (Xiong et al., 2024), another open-source innovation, expands the scope by serving as a generalist evaluator. Trained on diverse and detailed datasets, LLaVA-Critic provides robust scoring and preference learning, closely aligning with human and proprietary evaluations. These models mark significant progress in democratizing and enhancing multimodal evaluation tools.
Figure 7. LLM-as-a-Judge appears in two common forms in the agent. The left diagram is Agent-as-a-Juge, designing a complete agent to serve as an evaluator. The right diagram shows using LLM-as-a-Judge in the process of an Agent.
2.4.2. LLM-as-a-Judge for data
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of LLMs presents an unprecedented opportunity to automate the complicated process of data annotation by LLM-as-a-Judge. Most of the data need to be evaluated by LLM-as-a-Judge is generated by models, or large-scale crawled data. Language models first conduct supervised fine-tuning to imitate how to align with human instructions (Wang et al., 2023a; Taori et al., 2023). After that, reinforcement learning techniques have been explored to align language models with human preferences (Ouyang et al., 2022a; Ramamurthy et al., 2022). The most successful way is applying a RLHF framework (Ouyang et al., 2022a) via training a reward model on human feedback and using PPO (Schulman et al., 2017) to obtain the policy model for language generation. However, in practices, the PPO training paradigm is complex in coding and hyper-parameter tuning while it needs four models that are hard for training. This motivates us to explore simpler and more straightforward methods to align language models with human preferences. This involves how to use LLM-as-a-Judge to evaluate whether different responses are aligned with human preferences. For example,  (Yuan et al., 2023; Dong et al., 2023) use general LLM (ChatGPT) to get better alignment with human preferences. The Aplaca prompts (Taori et al., 2023) is used as sampling queries to different models generate responses. And these data was evaluated by LLM-as-a-Judge to obtain human preference scores (reward score) to train a new language model. Other works would like to use Supervised Fine-Tuning (SFT) model itself as evaluator, like generating better-aligned datasets for SFT including hindsight-modified prompts (Zhang et al., 2023a; Liu et al., 2023c) and principle-driven self-alignment (Sun et al., 2023b).
In addition, the lack of domain-specific model training data is a common phenomenon. In order to obtain annotated high-quality data, it is also very common to use LLM-as-a-Judge for the generation and evaluation of domain data. WizardMath (Luo et al., 2023a) would use its Instruction Reward Model (IRM) as Evaluator, aiming to judge the quality of the evolved instructions on three aspects: i) Definition, ii) Precision, and iti) Integrity. To produce the ranking list training data of IRM, for each instruction, ChatGPT and Wizard-E are used to generate 2-4 evolved instructions respectively. Then we leverage Wizard-E to rank the quality of those 4-8 instructions.
Recent research on evaluating multimodal data focuses on addressing vision-language misalignments in Multimodal Large Language Models (MLLMs), which often cause hallucinations—outputs inconsistent with visual or contextual evidence (Li et al., 2023a; Wang et al., 2023e; Cui et al., 2023). Techniques like Reinforcement Learning from Human Feedback (RLHF) and Factually Augmented RLHF have been employed to improve model alignment by incorporating structured ground-truth data and image captions, enhancing hallucination detection (Sun et al., 2023a). Benchmarks such as MLLM-as-a-Judge (Chen et al., 2024b) assess these models using tasks like scoring, pair comparison, and batch ranking, revealing limitations in alignment with human preferences. Persistent issues include biases (e.g., position, verbosity) and hallucinations, with even advanced models like GPT-4V displaying challenges. While pair comparison tasks align better with human judgment, scoring and batch ranking require significant improvements for reliable deployment. These findings emphasize the need for innovative frameworks and datasets to refine MLLM evaluation and alignment.
2.4.3. LLM-as-a-Judge for agent
There are two ways to apply LLM-as-a-Judge for an agent. One is to evaluate the entire process of the intelligent agent (Zhuge et al., 2024), and the other is to evaluate it at a specific stage in the agent framework process (Hao et al., 2023; Shinn et al., 2024). Both approaches are briefly illustrated in Figure 7. Using LLM as the brain of agent, an agentic system (Zhuge et al., 2024) could evaluate like a human, it would reduce the need for human involvement and eliminate the trade-off between thoroughness and effort. In addition, the agent (Shinn et al., 2024) can interact with the environment through language and receive feedback on actions through LLM to make decisions for the next action.
2.5. [Uncaptioned image] Quick Practice
Figure 8. Flowchart of Quick Practice
[…]
3. Improvement Strategy
When directly utilizing LLMs to conduct evaluation tasks such as scoring, selection, pairwise comparison or ranking, the inherent biases of LLMs like length bias, positional bias and concreteness bias(Park et al., 2024) will lead to poor evaluation results. Addressing these inherent biases and improving the overall evaluation performance of LLMs is a critical challenge for applying LLMs as evaluators. In this section, we introduce three improvement strategy aimed at enhancing the evaluation performance of LLM-as-a-judge: design strategy of evaluation prompts (in-context learning based), improvement strategy of LLMs’ evaluation capabilities (model based), and optimization strategy of final evaluation results (post-processing based). Our categorization is based on the formal definition of LLM-as-Evaluator in Section 2, focusing on enhancing the evaluation effectiveness by targeting three key phases of the process: the context $\mathcal{C}$, the abilities of LLMs themselves $\mathcal{P_{LLM}}$ and the post processing $\leftarrow$ to obtain the final results $\mathcal{E}$
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [Improvement
Strategy, fill=green!20 [Design Strategy of
Evaluation Prompts
(Sec. 3.1), text width=8em [Optimizing LLMs’
Understanding of
Evaluation Tasks, text width=6em [Few-shot prompting: FActScore (Min et al., 2023) / SALAD-Bench (Li et al., 2024b) / GPTScore (Fu et al., 2024)] [Evaluation steps decomposition: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024c) / SocREval (He et al., 2024b) / BSM (Saha et al., 2024)] [Evaluation criteria decomposition: HD-Eval (Liu et al., 2024b) / Hu and Gao et al. (Hu et al., 2024)] [Shuffling contents: Wang et al. (Wang et al., 2024a) / Auto-J (Li et al., 2023b) / JudgeLM (Zhu et al., 2023a) / PandaLM (Wang et al., 2023d)] [Conversion of evaluation tasks: Liu et al. (Liu et al., 2024c)] ] [Optimizing LLMs’
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024c) / LLM-EVAL (Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023) / FLEUR (Lee et al., 2024b)] ] ] [Improvement Strategy
of LLMs’ Abilities
(Sec. 3.2), text width=8em) [Fine-tuning via Meta
Evaluation Datasets, text width=7em [PandaLM (Wang et al., 2023d) / SALAD-Bench (Li et al., 2024b) / OffsetBias (Park et al., 2024) / JudgeLM (Zhu et al., 2023a) / CritiqueLLM (Ke et al., 2024)] ] [Iterative Optimization
Based on Feedbacks, text width=7em [INSTRUCTSCORE (Xu et al., 2023) / JADE (Zhang et al., 2023c)] ] ] [Optimization Strategy
of Final Results
(Sec. 3.3), text width=8em) [Integration of multiple
Evaluation Results, text width=7.5em [Summarize by multiple rounds: Sottana et al. (Sottana et al., 2023) / PsychoBench (Huang et al., 2023) / Auto-J (Li et al., 2023b)] [Vote by multiple LLMs: CPAD (Liu et al., 2023d) / Bai et al. (Bai et al., 2024b)] ] [Direct Optimization of
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024b) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024c)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]
Figure 9. Structure of Improvement Strategy.
3.1. Design Strategy of Evaluation Prompts
An evaluation prompt is an input to LLM evaluators, which is used to guide the LLMs to complete the required evaluation tasks. LLMs possess in-context learning ability to learn how to perform specified tasks through relevant examples or instructions provided in prompts without requiring weight updates or retraining(Brown, 2020). It indicates that the design strategy of evaluation prompts will significantly impact the effectiveness of LLM-as-a-judge. Therefore, how to optimize the design of evaluation prompts, including better methods to help LLMs understand the evaluation tasks and produce evaluation results, is the most direct and effective way to improve the evaluation performance of LLM-as-a-judge.
3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In optimization methods of prompting LLMs to better understand evaluation tasks, one of the most commonly used and effective approaches is few-shot prompting(Brown, 2020). By incorporating several high-quality evaluation examples into the evaluation prompts, LLM evaluators can effectively grasp the objectives, general processes and rough evaluation criteria of evaluation tasks. Many research works employ this prompt paradigm for evaluation, such as FActScore(Min et al., 2023), SALAD-Bench(Li et al., 2024b) and GPTScore(Fu et al., 2024).
In addition to providing hight-quality examples for LLMs to inference, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly including the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024c) use Chain-of-Thought(CoT)(Wei et al., 2022) to provide guidance for LLMs. SocREval(He et al., 2024b) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. proposes Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024b) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition and thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024a) analyzed and validated the impact of position bias on LLM-as-a-judge, and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.
To address the challenge of LLMs’ absolute scoring being less robust than relative comparing(Raina et al., 2024), some research works convert scoring tasks into pairwise comparison, thereby enhancing the reliability of evaluation results. Liu et al.(Liu et al., 2024c) transform the scoring evaluation to ranking evaluation and introduce Pairwise-Preference Search (PARIS), which employs LLMs to conduct pairwise comparisons locally and efficiently ranks candidate texts globally, making evaluation results more aligned with human preferences.
In summary, the design of prompts for better understanding evaluation tasks is a core method for optimizing LLMs’ in-contextual learning abilities. By refining evaluation task instructions and criteria in prompts or few-shot prompting with high-quality examples, the details of evaluation prompts can be enriched and the understanding of LLMs on evaluation tasks can be directly or indirectly enhanced. Additionally, targeted adjustments to prompts can address potential biases of LLMs such as position bias.
3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024c) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Chen, 2023) further codifies this form-filling paradigm, efficiently output evaluation results in JSON dictionary format and obtain multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
Apart from challenges in robustness, directly outputting evaluation results by LLMs also suffer from the lack of interpretability. The meaning of evaluation results from LLM evaluators is difficult to align consistently with instructions and metrics provided in prompts. To address the challenges, CLAIR(Chan et al., 2023) requires LLMs to output evaluation scores between 0-100 simultaneously with relevant reasons as explanations in JSON format, which enhancing the rationality and interpretability of the scores. FLEUR(Lee et al., 2024b) utilizes LLaVA to first provide quality scores for image captions and subsequently asks with ”Why? Tell me the reason.” for explanations with the images, captions and scores as inputs, offering a stepwise approach to provide interpretable scores.
In general, by constraining or guiding the output process and format of LLM evaluators within prompts, the robustness and rationality of evaluation results can be effectively improved through structured outputs. This also facilitates the automated post-processing of evaluation results in subsequent steps, thereby enhancing the overall stability of the evaluation pipeline.
3.2. Improvement Strategy of LLMs’ Abilities
The evaluation capabilities of LLMs is a reflection of their powerful general language understanding and generation abilities triggered by specific prompts. Methods for optimizing evaluation through prompt design, which focuses on LLMs’ in-contextual learning capabilities, require LLMs to fully comprehend the meaning of prompts and consistently follow the relevant evaluation instructions. However, even state-of-the-art LLMs like GPT4 face issues such as conceptual confusion(Hu et al., 2024), and smaller open-source LLMs which are easier to deploy as evaluators have even more limitations in their evaluation capabilities. Therefore, how to improve the evaluation capabilities of LLMs, including how to fine-tune LLMs through meta evaluation datasets and how to iteratively optimizing models based on feedback of evaluation results, is significant for improving the fundamental evaluation performance of LLM-as-a-judge.
3.2.1. Fine-tuning via Meta Evaluation Datasets
A straightforward approach to enhancing the evaluation capabilities of LLMs is to fine-tune them via meta evaluation datasets specifically constructed for evaluation tasks, which helps improve the LLMs’ understanding of specific evaluation prompts, boosts the evaluation performance, or addresses potential biases. The most critical step in this optimization strategy is the collection and construction of training data. A common method involves sampling evaluation questions from publicly available datasets, modifying them with certain templates, and supplementing the dataset with evaluation responses generated either manually or by powerful LLMs like GPT4. For instance, PandaLM(Wang et al., 2023d) samples inputs and instructions from Alpaca 52K(Taori et al., 2023) and generate responses using GPT-3.5 to construct training data, while SALAD-Bench(Li et al., 2024b) builds its training data from a subset of LMSYS-Chat(Zheng et al., 2023a) and Toxicchat(Lin et al., 2023).
To better align with the requirements of evaluation tasks, many research works further transform inputs and instructions sampled from public datasets to construct more targeted training data. OffsetBias(Park et al., 2024) aims to reduce biases of LLMs by using GPT4 to generate off-topic versions of the original inputs and then having GPT-3.5 respond to the new inputs to produce bad responses. By pairing good and bad responses as training data to fine-tune the LLMs as evaluators, the biases in LLMs are significantly reduced, including length bias, concreteness bias, knowledge bias and so on. JudgeLM(Zhu et al., 2023a) enhances LLMs’ evaluation capabilities by creating different types of training data through paradigms like reference support and reference drop. CritiqueLLM(Ke et al., 2024) proposes a multi-path prompting approach, combining pointwise-to-pairwise and referenced-to-reference-free prompting strategies to restructure referenced pointwise grading data into four types, which helps create Eval-Instruct to fine-tune LLMs, addressing shortcomings in pointwise grading and pairwise comparison.
In summary, constructing meta evaluation training data targeted at specific evaluation tasks and fine-tuning LLMs can directly adjust the model’s internal parameterized knowledge and language abilities. This is the most straightforward method to improve the evaluation performance of LLM evaluators and address potential biases.
3.2.2. Iterative Optimization Based on Feedback of Evaluation Results
Fine-tuning LLMs on meta evaluation datasets give them the ability to produce evaluations which are more aligned with human preferences. However, LLM-as-a-judge may still introduce biases during evaluation process in practice, which can impact the overall evaluation quality. A natural improvement strategy is to iteratively optimize the model based on feedback of evaluation results, which mainly comes from stronger models or directly from human evaluators’ correction of the evaluation results.
A typical example is INSTRUCTSCORE(Xu et al., 2023). To improve model performance and further benefit the final quality score calculation, this score framework collects failure modes of metric outputs, query GPT-4 on each failure mode to gather automatic feedback, and finally selects explanations most aligned with human preferences to iteratively fine-tune the LLaMA model. Unlike INSTRUCTSCORE which directly optimizes the model, the LLM evaluator in JADE(Zhang et al., 2023c) relies on human judges to correct LLMs’ evaluation results and updates the most frequently corrected samples into the example sets for few-shot prompting. JADE utilizes this relatively low-cost method to achieve iterative updates of the evaluation capabilities.
Since the feedback is more closely aligned with human preferences, LLM evaluators can dynamically align with human when optimizing evaluation capabilities based on this feedback, leading to better evaluation results. This feedback-based iterative optimization strategy address the problem of models’ imperfect generalization and improve the evaluation capabilities through dynamic updates.
3.3. Optimization Strategy of Final Results
Through the optimization based on in-context learning and the model’ own capabilities, LLMs have become fairly reliable evaluators which are capable of understanding evaluation task requirements and providing rational evaluation results. However, the inherent generation randomness within the black box of LLMs still introduces significant instability to the entire evaluation pipeline, affecting the overall evaluation quality. Therefore, optimization strategies during the post-processing stage from LLM evaluators’ outputs to final evaluation results are necessary. In this survey, these optimization strategies are categorized into three types: integration of multiple evaluation results, direct optimization of LLMs’ outputs, and conversion of evaluation tasks from pointwise evaluation to pairwise comparison.
3.3.1. Integration of Multiple Evaluation Results
Integrating multiple evaluation results for the same content to obtain the final result is a common strategy in various experiments and engineering pipelines, which can reduce the impacts of accidental factors and random errors. The most basic optimization strategy is to perform multiple runs of evaluation on the same content with different hyper-parameters and settings, and then summarize these results. For example the work of Sottana et al.(Sottana et al., 2023) reduces randomness in evaluations by averaging multiple scores of the same sample. Similarly, PsychoBench(Huang et al., 2023) takes the mean and standard deviation from ten independent runs. Auto-J(Li et al., 2023b) further amplifies the differences between evaluation rounds, which combine critiques with and without scenario criteria to obtain the final results.
In addition to integrating results from multiple rounds of evaluation, using multiple LLM evaluators to assess the contents simultaneously and the integrating the results is another effective method, which can reduce biases introduced by LLMs. For instance, CPAD(Liu et al., 2023d) utilizes ChatGLM-6B(Du et al., 2022), Ziya-13B(Zhang et al., 2022a) and ChatYuan-Large-v2(Zhang and Zhao, 2022) as evaluators to evaluate the contents and obtain the final results by voting. Bai et al.(Bai et al., 2024b) propose a novel evaluation method called decentralized peer review of LLMs, which utilizes LLMs that generate contents to evaluate each other’s generated contents and eventually integrate the results.
In summary, forming the final evaluation results by combining multiple rounds of evaluations or multiple LLM evaluators can reduce the random effects caused by accidental factors in a single round and reduce the potential biases of single LLM evaluator. This strategy significantly enhances the stability and reliability of the evaluation results.
3.3.2. Direct Optimization of LLMs’ Outputs
[…]
However, methods like score smoothing, which combine implicit logits and explicit outputs require the LLMs to be open-source, or to provide interfaces that allow access to token probabilities, which brings some limitations. Inspired by the work of Weng et al.(Weng et al., 2023) and Madaan et al.(Madaan et al., 2024), self-verification can be used to filter out the evaluation results without sufficient robustness. For example, TrueTeacher(Gekhman et al., 2023) applies self-verification in its evaluation of distilled data by asking the LLM evaluator for its certainty about the evaluation results after providing them, and retaining only those results that pass self-verification. Self-verification is suitable for all LLMs and require no complex computing and processing.
In summary, compared to integrating multiple evaluation results, directly optimizing the LLMs’ outputs to obtain the final results is faster and more low-cost, although the effectiveness still needs further validation. However, these two approaches are not mutually exclusive. Performing integration after direct optimization of LLMs’ output may lead to more stable evaluation results.
4. Evaluation of LLM Evaluators
Despite their impressive performance, LLMs exhibit several notable shortcomings, such as hallucinations (Tonmoy et al., 2024), biases (Gallegos et al., 2024), and a lack of robustness (Zhu et al., 2023d). When LLMs are employed as evaluators, these inherent issues can lead to suboptimal evaluation outcomes. Therefore, it is crucial to accurately and comprehensively assess the quality of LLM-as-a-judge and identify potential vulnerabilities. This section will review existing work on the evaluation of LLM-as-a-judge, focusing on three key areas: base metric (Section 4.1), bias (Section 4.2), and robustness (Section 4.3).
4.1. Basic Metric
The main objective of LLM-as-a-judge is to achieve alignment with human judges. Numerous studies approach this by considering the LLM evaluator as a virtual annotator and evaluating the extent of its agreement with human annotators. The percentage agreement metric represents the proportion of samples on which LLM and human annotators agree (Thakur et al., 2024).
$\text{Agreement}=\frac{\sum_{i\in\mathcal{D}}\mathbf{I}(\text{S}_{\text{llm}}=\text{S}_{\text{human}})}{\|\mathcal{D}\|}$
where $\mathcal{D}$ is the dataset, $S_{\text{llm}}$ and $S_{\text{human}}$ is the evaluation result of LLM evaluator and human judge respectively, which can be in the form of both score or rank. Additionally, widely used correlation metrics such as Cohen’s Kappa (Thakur et al., 2024) and Spearman’s correlation (Liu et al., 2024c; Bai et al., 2024b) are also employed to access agreement. Other works treat the LLM-as-a-judge task as a classification problem, where human annotations serve as the labels, and compute precision, recall, and F1 scores to evaluate the performance (Zhu et al., 2023b; Wang et al., 2023d).
Both of above metrics rely on the datasets with LLM-generated response and responding human judgements. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. In (Zheng et al., 2023d), MTBench and Chatbot Arena Conversations are proposed. The former has only 80 human-crafted queries, each with several LLMs’ responses and expert-level human annotation on pairwise comparison; the latter is a large collection of crowdsourced data, with more than 30K queries from real-world users and their vote on pairs of responses from different LLMs. FairEval (Wang et al., 2023b) is based on the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated labels between ChatGPT and Vicuna responses. PandaLM (Wang et al., 2023d) constructs a test set comprising 999 pairwise samples, with queries from 252 user-oriented instructions in (Wang et al., 2022). LLMEval2 (Zhang et al., 2023d) is much larger than the previous two, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Shepherd (Wang et al., 2023c) collects 352 samples from multiple sources for its critique model as a test set to evaluate the quality of the critiques. Table 1 shows the benchmarks and their statistics.
Current meta-evaluation primarily focuses on LLM-as-a-judge for models, while there is a lack of sufficient meta-evaluation when these LLM evaluators are used for automatically annotating large-scale datasets (Section 2.4.2). We advocate for more rigorous accessment of the alignment between LLM-as-a-judge and human judgment when they are employed for large-scale data annotation. Additionally, it is also crucial to assess the potential bias and robustness, which will be discussed in the following sections.
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023d) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023d) 2023 30k Pairwise 3 FairEval (Wang et al., 2023b) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023d) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023d) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023c) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12
Table 1. Benchmark for meta-evaluation of LLM-judge.
4.2. Bias
In this section, we systematically review various types of biases in the LLM-as-a-judge context, including their definitions, relevant metrics, and datasets that can be used for evaluation.
Position Bias is the tendency of LLM evaluators to favor responses in certain positions within the prompt (Wang et al., 2024a; Thakur et al., 2024; Shi et al., 2024; Ye et al., 2024). This bias may have detrimental effects, as Vicuna-13B could outperform ChatGPT when evaluated by ChatGPT, simply by positioning the response of Vicuna-13B in the second place (Wang et al., 2024a). To measure this bias, (Shi et al., 2024) proposed two metrics: Position Consistency, which quantifies how frequently a judge model selects the same response after changing their positions, and Preference Fairness, which measures the extent to which judge models favor response in certain positions. (Wang et al., 2024a) introduced a metric Conflict Rate to measure the percent of disagreement after change the position of two candidate responses. Their analytical experiments reveal that the degree of positional bias fluctuates depending on the disparity in response quality and the preferred position varies with different LLMs. For instance, GPT-4 tends to favor the first position, while ChatGPT shows a preference for the second position.
Length Bias refers to the tendency to favor responses of a particular length, such as a preference for more verbose responses which is also known as verbosity bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023d; Ye et al., 2024). Length bias can be revealed by rephrasing one of the original response into a more verbose one (Ye et al., 2024; Zheng et al., 2023d). Even though these expansions do not introduce new information, there is still concern regarding changes to the original response in terms of perplexity, fluency, or style. Alternatively, (Saito et al., 2023) investigated this bias by comparing multiple sampled responses and revealed a statistical tendency towards longer answers. However, ensuring the comparable quality of multiple samples remains a challenging problem.
Self-Enhancement Bias describe the phenomenon that LLM evaluators may prefer response generated by themselves (Zheng et al., 2023d; Ye et al., 2024). Considering the significant self-enhancement bias, (Ye et al., 2024) suggested that we should avoid using the same model as the evaluator. This is only a stopgap, as we may not use the optimal evaluator when evaluating the most advanced LLMs.
Other Bias. Diversity Bias refers to bias against certain demographic groups (Ye et al., 2024), including certain genders (Chen et al., 2024a), race, and sexual orientation (Kumar et al., 2024b). (Chen et al., 2024a) revealed evaluators’ tendency toward visually appealing content, regardless of its actual validity, such as the text with emoji. Concreteness bias reflects that LLM evaluators favor responses with specific details, including citation of authoritative sources, numerical values and complex terminologies, which is called authority bias (Park et al., 2024) or citation bias (Chen et al., 2024a; Ye et al., 2024). Furthermore, LLM evaluators may favor response with certain emotional tones, such as cheerful, sad, angry, and fearful, which is defined as seyiment bias (Ye et al., 2024; Li and Sinnamon, 2023).
To advance the development of LLM-as-a-Judge systems, future efforts should address two key challenges: (i) Need for Systematic Benchmark. Due to the diversity of biases, it is crucial to propose a systematic benchmark to evaluate the extent of various biases. As shown in Table 1, (Park et al., 2024) proposed EVALBIASBENCH as a test set to measure six types of bias. (Ye et al., 2024) is dedicated to proposing a unified bias testing process, including automated perturbation and a unified metric. They constructed a bias quantification framework CALM, which covers 12 types of bias. Despite these efforts, there is still no systematic benchmark and dataset that includes all types of biases. (ii) Challenges of Controlled Study. When conducting an investigation into a certain type of bias, it is challenging to isolate the specific direction of interest from other biases and quality-related characteristics. For instance, in the case of position bias, lengthening the response could potentially alter the style, fluency, and coherence, or even introduce new biases such as self-enhancement bias. Additionally, the tendency for GPT-4 to favor its own responses over those of GPT-3.5 can be interpreted as either self-enhancement bias or a proper tendency towards higher quality text. Therefore, it is essential for analytical work to carefully control for these variances.
4.3. Adversarial Robustness
Adversarial robustness refers to the ability of a model to withstand deliberate attempts to manipulate the scores through carefully crafted inputs. Unlike bias evaluations (Section 4.2) which mainly focus on naturally occurring samples, adversarial robustness involves samples intentionally crafted to manipulate scoring, such as inserting phrases that artificially enhance scores. Robustness is crucial because insufficient robustness allows trivial manipulations to deceive the evaluators and to undermine the evaluation of text quality. Ensuring robust evaluators is essential for maintaining accurate and reliable assessments, particularly in high-stakes applications.
(Raina et al., 2024) constructed a surrogate model from the black-box LLM-evaluator and the learn a adversarial attack phrases based on it. The evaluation score can be drastically inflated by universally inserting the learned attack phrases without improving the text quality. Furthermore, (Zheng et al., 2024) demonstrated that even a ”null model” that outputs a constant response irrelevant to input instructions can achieve high win rates for various LLM-as-a-judge methods. (Ye et al., 2024; Koo et al., 2023) proposed to increase the evaluation score by adding the majority opinions, such as “90% believe this is better”. (Ye et al., 2024; Koo et al., 2023) evaluated robustness against meaningless statement in the System Prompt, e.g., “”Assistant A loves eating pasta”. These works revealed that LLM-as-a-judge are still insufficiently robust against interference irrelevant to text quality. Defensive measures like the perplexity score (Raina et al., 2024; Jain et al., 2023) can only detect limited types of adversarial examples. Therefore, constructing more robust LLM-as-a-judge is a crucial research direction for the future.
5. Meta-evaluation Benchmark
In Section 3, We have introduced improvement strategies adopted by researchers in existing LLM-as-a-judge works to improve the evaluation capabilities of LLMs. Although numerous works have proposed meta-evaluation benchmarks to assess the performance of LLMs in evaluation tasks, as shown in Table 1, there is still a lack of meta-evaluation on whether these improvement strategies effectively optimize the LLM evaluators and which dimensions of evaluation performance are being enhanced. It is possible that some improvement strategies fail to enhance the LLM evaluators’ performance or mitigate biases in practical use, leading to computing waste. In this section, based on benchmarks mentioned in Section 4, we organized a robust and scalable meta-evaluation pipeline and benchmark, and conducted a simple meta-evaluation experiment on the improvement strategies summarized in Section 3, examining their effectiveness from the perspectives of biases and agreement with human evaluation.
5.1. Experiment Settings
5.1.1. Evaluation Dimensions and Benchmarks
The most direct meta-evaluation metric for reflecting the quality of automatic evaluation is the alignment with human evaluation. We use LLMEval2 (Zhang et al., 2023d) to assess the alignment of LLM-as-a-judge with human evaluations. LLMEval2 is the largest and most diverse evaluation benchmark for LLM-as-a-judge to date, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Each sample consists of a question, a pair of candidate responses, and a human label indicating the preferred response.
Bias is also a crucial dimension for assessing the quality of LLM-as-a-judge’s evaluation results. We use EVALBIASBENCH(Park et al., 2024) to measure six types of biases in LLMs, including length bias, concreteness bias, empty reference bias, content continuation bias, nested instruction bias, and familiar knowledge bias. EVALBIASBENCH consists of 80 samples, each containing a question, a pair of candidate responses, and a label indicating the correct response without bias influence. In addition to the six types of biases, we also evaluated position bias. The meta-evaluation samples for position bias are the paired samples constructed by swapping the position of candidate responses within prompts in samples of LLMEval2 and EVALBIASBENCH.
5.1.2. Evaluation Metrics
For the alignment with human evaluation, we use the Percentage Agreement Metric for evaluation(Thakur et al., 2024), as shown in Section 4.1. For biases, we use the Accuracy for evaluation, which represents the proportion of samples on which LLMs select the correct candidate response annotated in EVALBIASBENCH.
For position bias, we use Position Consistency as metric, which quantifies how frequently a judge model selects the same response after changing their positions. Formally, given $N$ samples $\{(q_{i},r1_{i},r2_{i})\}^{N}_{i=1}$, for each sample $(q_{i},r1_{i},r2_{i})$, we query the LLM evaluator with two prompts $P(q_{i},r1_{i},r2_{i})$ and $P(q_{i},r2_{i},r1_{i})$, and obtain corresponding two evaluation results $S_{i}^{r12}$ and $S_{i}^{r21}$. Each $S_{i}$ is $r1_{i}$, $r2_{i}$ or ”TIE”. Then we calculate the position consistency as follows:
Position $\text{Position Consistency}=\frac{\sum_{i=1}^{N}\mathbb{I}(S_{i}^{r12}=S_{i}^{r21})}{N}$
where $\mathbb{I}(\cdot)$ is the indicator function.
5.1.3. Target LLMs and Strategies
For LLMs, we selected three LLMs commonly used in automatic evaluation as evaluators, including closed-source LLMs GPT-4, GPT-3.5, and open-source LLMs LLaMA3. For improvement strategies, we selected Providing Evaluations with Explanations, Self Validation, Summarize by Multiple Rounds, and Vote by Multiple LLMs, since these strategies are all very straightforward and relatively uniform in many works. We adopt GPT-3.5 as the base evaluator when conducting meta-evaluation for these improvement strategies.
5.1.4. Model Configuration
For the closed-source LLMs, we use the official API provided by OpenAI to interact with LLMs. The model versions we selected are GPT-4-turbo and GPT-3.5-turbo, specifically referencing gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125 respectively [^2] https://platform.openai.com/docs/models. For the open-source LLM, we adopt Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct provided by meta-llama on Hugging Face, deployed on an Ubuntu machine equipped with a 40GB NVIDIA A100 GPU. To stabilize the evaluation results of LLMs, we set the hyper parameter temperature to 0 to reduce the impact of randomness in LLMs’ output. For Summarize by Multiple Rounds, we conduct 5 rounds for each sample.
5.2. Experiment Results and Analysis
Model Biases Alignment Position Length Concre- Empty Content Nested Familiar with Human teness Reference Continuation Instruction Knowledge (n=5106) (n=2633) (n=34) (n=28) (n=26) (n=24) (n=24) (n=24) GPT-4-turbo 61.57 80.49 91.18 89.29 65.38 95.83 70.83 100.0 LLaMA3-8B-Instruct 50.72 38.85 20.59 57.14 65.38 75.00 45.83 54.17 GPT-3.5-turbo 54.72 68.78 20.59 64.29 23.08 91.67 58.33 54.17 GPT-3.5-turbo - w/ explanation 52.47 48.97 35.29 60.71 38.46 91.67 41.67 50.00 - w/ self-validation 54.86 69.31 23.53 60.71 23.08 91.67 41.67 50.00 - w/ multi rounds 54.68 70 11 26.47 67.86 23.08 95.83 54.17 50.00 Multi LLMs 57.66 32.28 26.47 64.28 46.15 87.50 66.67 62.50
Table 2. The meta-evaluation results for LLMs and improvement strategies based on GPT-3.5-turbo. All the values are percentages.
The experiment results are shown in Table 2. Comparing the evaluation performance of different LLMs, we found significant differences between them. GPT-4 outperforms GPT-3.5 and LLaMA3-8B-Instruct with a large margin across all meta-evaluation dimensions and shows fewer biases. Therefore, when economically feasible, using GPT-4 for automatic evaluation may obtain objective results with minimal biases. LLaMA3-8B-Instruct and GPT-3.5 have similar metrics across most meta-evaluation dimensions. However, LLaMA3-8B-Instruct performs poorly in position bias, while GPT-3.5 struggles in empty reference bias. Since LLaMA3-8B-Instruct is open-source, it may also achieve good evaluation performance as a relatively lightweight automatic evaluator after specific debiasing fine-tuning.
Comparing the performance of different improvement strategies, it reveals that no all modifications effectively improve LLM-as-a-judge’s evaluation outcomes. Providing with Explanation (- w/ explanation) provides interpretability by offering reasons alongside evaluation scores or selections, which aids in logical backtracking during human review. However, in terms of evaluation performance and bias mitigation, it generally has a negative impact. This performance decline is speculated to be caused by deeper biases introduced by self-explanation. Self Validation (- w/ self-validation) shows minimal effectiveness, likely due to the LLMs’ overconfidence, which may limit its re-evaluation efforts during self-validation. We will further discuss this limitation in Section 7.1.
Summarize by Multiple Rounds (- w/ multi rounds) is a strategy with clear benefits, showing improvements across multiple dimensions. It suggests that repeated evaluations help reduce the impact of randomness in LLMs, thereby addressing bias issues. Vote by Multiple LLMs (Multi LLMs) shows a noticeable decline in performance of position bias, but with either slight improvements or minimal differences in other dimensions, when compared to GPT-3.5-turbo. This trend is influenced by the comprehensive performance of GPT-4-turbo, GPT-3.5-turbo, and LLaMA3-8B-Instruct. GPT-4-turbo generally aids in correction, but it is significantly impacted in position bias, where LLaMA3-8B-Instruct underperforms. Similarly, GPT-3.5-turbo’s weak performance in handling empty reference bias notably reduces the corresponding metric. This indicates that when multiple LLMs are used for joint evaluation, the differences in their capabilities need to be carefully considered.
In summary, due to the inherent capabilities and potential risks of LLMs, common improvement strategies for LLM-as-a-judge are not fully effective in improving the evaluation performance or mitigating biases. The limitations and challenges will be further discussed in Section 7.
6. Application
LLMs’ ability as evaluators has gained widespread recognition in specialized fields, especially in complex, qualitative areas like legal texts, mathematical reasoning, and scientific research. This section reviews recent developments in LLM-as-a-judge applications across finance, law, science, and other industries, investigating how domain knowledge and LLM evaluator can further expand their impact in critical areas.
6.1. Machine Learning
6.1.1. NLP
LLMs have been successfully employed as evaluators in several NLP tasks, including sentiment analysis, machine translation, and text summarization. In sentiment analysis, numerous biases influencing LLM-based judgments have been identified, prompting the creation of automated frameworks to systematically quantify these biases. Despite progress, issues like self-preference and verbosity biases continue to exist (Ye et al., 2024). For translation, studies have shown that the effectiveness of LLM evaluators depends heavily on their English training, creating limitations in assessing cultural and factual accuracy in non-English contexts (Son et al., 2024). In text summarization, comprehensive guides have been developed to assist in evaluating LLM-generated content, stressing the importance of new metrics to better capture semantic qualities and minimize hallucinations (van Schaik and Pugh, 2024).
6.1.2. Multi-Modal AI Applications
In the field of multi-modal AI, benchmarks have been created to assess LLM-based systems that function across text and vision modalities. These benchmarks have enabled the evaluation of tasks such as image captioning and mathematical reasoning, where LLMs aligned with human preferences in pairwise comparisons but performed poorly in scoring and batch ranking (Chen et al., 2024b). For Chinese multi-modal alignment, benchmarks have identified challenges in coherence and reasoning, leading to the proposal of a calibrated evaluation model that achieves greater consistency than existing systems (Wu et al., 2024). Furthermore, advancements in multi-modal and multi-agent systems have been reviewed, emphasizing collaboration mechanisms to improve rationality and minimize biases (Jiang et al., 2024).
6.2. Other specific domains
6.2.1. Finance
LLMs have demonstrated significant potential in the finance domain, particularly in tasks such as forecasting, anomaly detection, and personalized text generation (Zhao et al., 2024), thereby driving an increasing demand for LLM evaluators.
In the context of LLM-as-a-judge applications within finance, expert knowledge is crucial for domain-specific evaluations. Current research can be divided into two areas: one focuses on designing LLM-based evaluators that leverage expert knowledge for specific tasks. For instance, Brief et al. (2024) conducted a case study on multi-task fine-tuning in finance to enhance LLM performance (Brief et al., 2024), while Yu et al. (2024) introduced FinCon, a multi-agent system that uses conceptual verbal reinforcement to improve financial decision-making (Yu et al., 2024). The second area of research aims to provide benchmarks to evaluate and enhance LLMs’ understanding of domain-specific knowledge. These benchmarks include user-feedback-based UCFE (Yang et al., 2024), IndoCareer—a dataset of professional exam questions (Koto, 2024), and AI-generated domain-specific evaluation sets (Raju et al., 2024).
Additionally, the concept of LLM-as-a-judge shows promising applications in credit scoring (Yoon, 2023; Babaei and Giudici, 2024) and Environmental, Social, and Governance (ESG) scoring (Zhao et al., 2024). This work remains in its early stages, necessitating further exploration to refine evaluation methods and expand applications in the finance domain.
6.2.2. Law
LLMs have shown growing capabilities in providing professional advice in specialized fields such as legal consultation, particularly excelling in tasks like text summarization and legal reasoning. However, compared to other fields, the legal sector is more concerned about potential biases and factual inaccuracies within LLMs. Similar to the finance domain, existing research in law can be divided into two main categories.
The first category focuses on developing LLM evaluators specifically for legal applications by addressing professional limitations or designing evaluators themselves. For example, Cheong et al. (2024) propose a four-dimensional framework for constructing responsible LLMs for legal advice, emphasizing (a) user attributes and behaviors, (b) the nature of queries, (c) AI capabilities, and (d) social impacts (Cheong et al., 2024). Ryu et al. (2023) developed Eval-RAG, a retrieval-augmented generator (RAG)-based evaluator that assesses the validity of LLM-generated legal texts. Testing on a Korean legal question-answering task, they found that combining Eval-RAG with traditional LLM evaluation methods aligns more closely with human expert evaluations (Ryu et al., 2023).
The second category of research involves creating benchmarks for evaluating LLM applicability in legal scenarios. Examples include multi-domain evaluation sets, such as the IndoCareer dataset for professional exams in Indonesia (Koto, 2024) and LegalBench, a collaboratively built benchmark for assessing legal reasoning capabilities in LLMs across multiple domains and languages (Guha et al., 2024). These benchmarks are often language-specific like LexEval for Chinese legal texts (Li et al., 2024a) and Eval-RAG for Korean  (Ryu et al., 2023). Other benchmarks target specific attributes, such as ethics (Zhang et al., 2024) and harmfulness (Andriushchenko et al., 2024).
6.2.3. AI for Science
LLMs have demonstrated notable potential in scientific fields, especially in areas like medical question-answering and mathematical reasoning, where they serve as evaluators to improve accuracy and consistency. In medical applications, studies by Brake et al. (2024) and Krolik et al. (2024) showed that models like LLaMA2 can assess clinical notes and Q&A responses with a level of accuracy approaching that of human experts (Brake and Schaaf, 2024; Krolik et al., 2024). This approach leverages prompt engineering to embed expert knowledge, enabling LLMs to handle complex, nuanced information, which provides a reliable first-line assessment that lessens the load on human experts.
In mathematical reasoning, reinforcement learning (RL) and cooperative reasoning methods further enhance LLM’s capability as an evaluator, especially for theory-proofing works (Lu et al., 2023). For example, WizardMath was introduced by employing RL through step-by-step feedback to refine reasoning in mathematical tasks (Luo et al., 2023b). Zhu et al. (2023) proposed a Cooperative Reasoning (CoRe) framework that combines generation and verification to mimic human-like dual-process reasoning, enhancing the model’s problem-solving accuracy (Zhu et al., 2023c). Additionally, Lu et al. (2023) developed MathVista, a benchmark for evaluating mathematical reasoning in visual contexts, which assesses LLMs like GPT-4V on tasks involving mathematical reasoning with visual components (Lu et al., 2023). These methods highlight the value of combining RL, cooperative reasoning, and prompt engineering in improving LLMs’ evaluative and reasoning skills across mathematical reasoning.
6.2.4. Others
LLMs have been employed as evaluators across various fields to enhance efficiency and consistency. In software engineering, a method was proposed for using LLMs to evaluate bug report summarizations, demonstrating high accuracy in assessing correctness and completeness, even surpassing human evaluators who experienced fatigue. This approach offers a scalable solution for evaluation (Kumar et al., 2024a). In education, automated essay scoring and revising were explored using open-source LLMs, achieving performance comparable to traditional deep-learning models. Techniques such as few-shot learning and prompt tuning improved scoring accuracy, while revisions effectively enhanced essay quality without compromising original meaning (Song et al., 2024). In content moderation, an LLM-based approach was developed to identify rule violations on platforms like Reddit, achieving high true-negative rates but encountering challenges with complex rule interpretation, emphasizing the necessity of human oversight for nuanced cases (Kolla et al., 2024). These applications of LLMs as evaluators highlight their growing potential in diverse sectors, emphasizing the need for integrating domain-specific knowledge and refining methodologies.
Moreover, LLMs as evaluators demonstrate significant advantages in qualitative assessments that are difficult to quantify, such as evaluating service quality, analyzing user experience feedback, and assessing creative content like art or literature reviews. LLMs’ capability to understand and generate nuanced language makes them well-suited for subjective evaluation tasks traditionally requiring human judgment. Future research will increasingly focus on these areas, exploring how LLMs as judges can enhance assessment accuracy and consistency where traditional quantitative methods fall short.
7. Challenges
In this chapter, we explore the key challenges that arise when utilizing LLMs for evaluation tasks, particularly in the context of LLM-as-a-Judge. Despite their growing capabilities, LLMs still face significant issues related to reliability, robustness, and their backbone models’ limitations. Understanding these challenges is crucial for advancing the use of LLMs in a fair, consistent, and reliable manner. We address these concerns under three main themes: reliability, robustness, and the need for more powerful backbone models.
7.1. Reliability
Evaluating the reliability of LLMs when used as judges reveals several pressing challenges. Both human and LLM judges exhibit biases, which raises concerns regarding the consistency and fairness of their evaluations. Specifically, human judges are also found to have inherent bias (Zheng et al., 2023b; Wu and Aji, 2023) and may not even provide reliable answers (Clark et al., 2021; Hämäläinen et al., 2023). As an alternative to humans, LLM evaluations are also found to have certain biases, and the annotation results require more evaluation (Pangakis et al., 2023), as we discussed in § 4. The bias of LLM-as-a-judge is more due to the fact that LLM is a probabilistic model, as we have defined in § 4. Moreover, Reinforcement Learning with Human Feedback (RLHF) improves LLM performance by aligning them with human preferences. However, ensuring models trained with RLHF produce robust and consistent outputs remains an ongoing challenge.
In this section, to better understand reliability, we discuss the reliability issues that arise from biases, overconfidence, and challenges in generalization.
Overconfidence
Instruction-tuned LLMs have been demonstrated to possess the issue of overconfidence, which means they tend to offer overly favorable scores when evaluating their own responses (Tian et al., 2023a). The overconfidence is also highly likely to exist in the scenario of LLM-as-a-judge, which is also engaged in evaluating the responses generated by LLMs. Consequently, when LLM-as-a-judge is utilized with the latest LLMs, which are typically instruction-tuned, the existence and impact of overconfidence need to be meticulously examined.
Fairness and Generalization
Another significant aspect of reliability is fairness and generalization. Evaluations by LLM-as-a-judge can exhibit considerable inconsistency depending on the context. This is why prompt-based methods are often used to improve LLM-as-a-judge performance. However, challenges related to fairness and generalization may arise due to the sensitivity of prompt engineering.
For example, the order of the examples in the context can significantly affect the model’s output, leading to unfair evaluations if the examples are poorly arranged. Moreover, LLMs struggle to handle long context windows effectively, often showing degraded performance or prioritizing later examples in the sequence. These issues raise concerns about fairness and generalization in LLM-based evaluations.
7.2. Robustness
Despite LLM’s superior power, it is found prone to adversarial attacks (Shen et al., 2023; Jiang et al., 2023; Zou et al., 2023), under which LLMs can be induced to generate harmful content. While existing works on LLM attacks mainly focus on NLG tasks, more attacks on LLM-as-a-judge are relatively under-explored (Chen et al., 2024a). This means that we will face some robustness challenges when using LLM-as-a-Judge, and these risks are unknown.
Addressing these robustness challenges requires a deeper understanding of the specific vulnerabilities associated with LLM-as-a-Judge tasks. Unlike traditional adversarial attacks on natural language generation (NLG), where the goal is often to mislead the model into generating harmful or incorrect outputs, attacks on LLM-as-a-Judge aim to exploit biases, inconsistencies, or loopholes in the model’s decision-making processes. For instance, subtle manipulations in input phrasing or context framing could potentially lead to significant deviations in judgments, raising concerns about reliability in high-stakes applications.
Currently, we have some methods to defend against such attacks to maintain robustness. These approaches mainly involve post-processing techniques, such as response filtering and consistency checks, which are essential for improving evaluation quality. However, these techniques still face significant challenges. One major issue is self-consistency, as LLMs often produce inconsistent outputs when evaluating the same input multiple times. Another challenge is random scoring, where the model assigns arbitrary or overly positive scores that fail to accurately reflect the true quality of the generated outputs. Such limitations undermine the reliability and robustness of these defense mechanisms.
7.3. Powerful Backone Model
Although LLMs show superior performance in text-based evaluation, the field lacks robust multi-modal models to effectively serve as reliable judges for multi-modal content. Current multi-modal LLMs, such as GPT-4 Vision, still struggle with complex reasoning across different modalities. This limitation poses a challenge to achieving reliable evaluations on multi-modal assessment tasks. Even in many cases, our LLM cannot complete high-quality evaluation content due to insufficient powerful instruction-following ability and reasoning ability for evaluating text content.
8. Future Work
The rapid advancements in LLMs open new avenues for innovation while posing significant challenges for their application in critical domains. This section outlines potential directions to enhance the reliability and versatility of LLMs, focusing on their role as evaluators, annotators, and multi-modal evaluators to address pressing needs in research and industry.
8.1. More Reliable LLM-as-a-Judge
As highlighted in our Formulation (§ 2) and Strategy (§ 3), LLMs are probabilistic models that require extensive research and optimization to enhance their reliability as judges. Although current methods have improved the reliability of LLM-as-a-Judge, many challenges, including adaptability and robustness, remain unresolved. To enable probabilistic models to deliver evaluations closely aligned with real-world scenarios, future research should prioritize refining and implementing LLM-as-a-Judge across the evaluation pipeline.
There is considerable potential for improving reliability in various aspects, including In-Context Learning, model selection, post-processing techniques, and the overall evaluation framework for LLM-as-a-Judge. These efforts should prioritize not only enhancing the reliability of assessments but also developing methodologies to systematically evaluate and validate the robustness of these assessments. Furthermore, the establishment of comprehensive evaluation benchmarks and interpretable analytical tools will be crucial for assessing and improving the reliability of LLM evaluators.
Finally, the uncertain and evolving nature of robustness risks underscores the necessity of proactive mitigation strategies. These strategies should include the development of adversarial training techniques tailored to judgment tasks, the integration of robust uncertainty quantification methods, and the implementation of human-in-the-loop systems to oversee critical decisions. By addressing these challenges, we can build more resilient and dependable systems capable of maintaining high levels of reliability even under adversarial conditions.
8.2. LLM-as-a-Judge for Data Annotation
Despite its wide applications, data annotation poses significant challenges for current machine learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 (OpenAI, 2023a), Gemini (Google, 2023), and LLaMA-2 (Touvron et al., 2023c) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2022b), ensure consistency across large volumes of data, and adapt through fine-tuning or prompting for specific domains (Song et al., 2023), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP.
Whether in the field of scientific research or industry, we are all still suffering from insufficient target data and domain-specific data, or situations where the data quality is not high enough. Assuming that LLM-as-a-judge can achieve stable performance and be fair and reliable, we can use LLM to annotate data in scenarios where data is insufficient to expand the data. In scenarios with low data quality, we can assess the data quality through LLM, and label the quality tags to achieve the goal of selecting high-quality data. Currently, we have not been able to experimentally rely solely on LLM for a reliable evaluation of various different scenarios of data; most of the time, we still rely on human annotation to ensure professionalism and reliability. LLM-as-a-judge often needs to learn from human annotations in order to perform certain labeling tasks.
8.3. MLLM-as-a-Judge
AI systems are rapidly evolving into highly multifunctional entities. For example, whereas in the past we had special-purpose solutions for different language processing tasks (e.g., sentiment analysis, parsing, dialogue), LLMs are competent at all these tasks using a single set of weights (Srivastava et al., 2022). Unified systems are also being built across data modalities: instead of using a different architecture for processing images versus text, recent models, such as GPT4-V (OpenAI, 2023b), Gemini (Google, 2023), and LLaVA (Liu et al., 2023b), handle both modalities In brief, AI systems are becoming more uniform in their structures and functions. This also applies to LLM-as-a-Judge.
The future of evaluation lies in developing robust multi-modal evaluators that can handle complex, multi-modal content, such as text, images, and video. Current multi-modal LLMs, while promising, lack the reasoning depth and reliability of their text-based counterparts. Future research should focus on:
A practical multi-modal evaluator will not only advance research but also open new possibilities for tasks such as multi-modal content moderation and knowledge extraction.
8.4. More LLM-as-a-Judge Benchmarks
The development of more comprehensive and diverse benchmarks is also critical for advancing the reliability and applicability of LLM-as-a-Judge systems. Future efforts could focus on creating high-quality, large-scale datasets that encompass a wide range of scenarios, including domain-specific applications, multi-modal content, and real-world complexities. Additionally, benchmarks should integrate more detailed and fine-grained evaluation metrics. These improvements will not only provide a more holistic understanding of LLM performance but also guide the development of methodologies to enhance their capabilities. By establishing rigorous standards and datasets akin to ImageNet (Deng et al., 2009) in scale and impact, the LLM-as-a-Judge field can achieve deeper insights and foster greater innovation.